Need it answered step by step.
expect to receive on between the the firms deci- more of its -4 Suppose that five years ago Cisco Systems sold à 15.year böhd issue that Had a Bund vatddt $1,000 par value and a 7 percent coupon rate. Interest is paid semiannually a. If the going interest rate has risen to 10 percent, at what price would the bonds be selling today? Suppose that the interest rate remained at 10 percent for the next 10 years. What would happen to the price of Ciscos bonds over time? b.
Answer #1

a: Price = $813.07

Using financial calculator; Inputs

FV = 1000; N= 10*2 = 20 ; PMT =7%*1000/2 = 35 ;

Find PV as 813.07

In excel we can find thePV using PV function as per formula in the image

AutoSave Book1 - Excel 《Product Activation Failed) Sign in File Insert Page Layout Formulas Data Review View Help Tell me what you want to do 수 Share Home Calibri Wrap Text Currency 田FF Paste u . 녀 . 2.A 호 들經垣@werge & Center-5 . % , Conditional Fornat as Cell Insert Delete Fornat FormattingTable Styles- 、. Sort & Find & : Filter-Select- Clipboard Alignment Number Styles Cells Editing B2 x -PV(10%/2,20,35,1000) ($813.07) 10 12 13 14 15 16 17 Sheet1 Ready + 100 ENG 11:16 PM

b: If the YTM remained at 10%, the price of the bond will slowly reach its par value by the end of 10 years.